There are many comets in our solar system. A comet is located at a distance rc = 16 AU from the sun moving at a speed vc = 0.26vE where vE is the average speed of Earth orbital motion. At that particular instant the position vector rc makes an angle ϕ = 68o with velocity vector vc .

a) Determine eccentricity ε of the orbit.

b) If the semi-major axis of the orbit a = 13.7 AU, write the equation of the orbit for the comet.

c) Determine the period of the comet. Can you identify the most probable name for the comet?


1
Expert's answer
2021-01-27T16:49:29-0500

1)

"r=a(1-ecos (180°-\phi))," "\u21d2"

"e=\frac{1-\frac ra}{cos(180°-\phi)},"

"e=\frac{1-\frac {16}{13.7}}{cos(112°)}\approx0.4482."

2)

"r=\frac{p}{1+ecos\theta},"

"p=a(1-e^2)=10.9484,"

"r=\frac{10.9484}{1+0.4482cos\theta}."

3)

"T=a^{\frac 32}=50.7" years.

Comet 166P/NEAT.